OpenStudy (anonymous):

Will FAN and MEDAL ONLY need HELP with PART C In a certain town the temperature, x in degrees Celsius on a certain day is described by two statements: If 3 times the temperature is increased by 2, the temperature is still less than 14°C. Twice the temperature minus 7 is greater than -11°C. Part A: Create a compound inequality to represent the temperature range. (3 points) Part B: Can the temperature in this town be 5°C? Justify your answer by solving the inequalities in Part A. (3 points) Part C: The average temperature in another town is 3°C, but the actual temperature is

OpenStudy (anonymous):

37.4 farenhient

OpenStudy (anonymous):

37.4 i think

OpenStudy (anonymous):

how @coltjoco

OpenStudy (anonymous):

can you explain

OpenStudy (anonymous):

use a Celsius to Fahrenheit conversion

OpenStudy (anonymous):

Part C: The average temperature in another town is 3°C, but the actual temperature is within 4°C of the average. Write and solve an inequality to find the range of temperature in this town. (4 points) part of Part C got cut off here is the rest

OpenStudy (triciaal):

3x +2 <14 2x -7 >-11 x<4 x>-2 -2<x<4 no 5 > 4 when x = 3 (3) + 2 <14 2(3) - 7>-11 -1<x<10

OpenStudy (anonymous):

@triciaal can you explain the second part?

OpenStudy (triciaal):

without the rest of the question for the other town what i did was use the x = 3 to see what it would be in the first town.

OpenStudy (anonymous):

sorry still don't get it

OpenStudy (triciaal):

i guess it didn't prove anything except that 3 is in the range

OpenStudy (triciaal):

with the rest of the information range is + 4 /-4 3 + 4 =7 3 - 4 = -1 temp in this town is between -1 and 7